OpenStudy (anonymous):

Which of the following constants can be added to x2 - x to form a perfect square trinomial? a)1/2 b)1/4 c) 1

OpenStudy (cwrw238):

divide coefficient of x by 2 and square it

OpenStudy (cwrw238):

coefficient of x = -1 ( -x)

OpenStudy (anonymous):

what is a perfect square trinomial'

OpenStudy (cwrw238):

a trinomial whose factors are equal eg (x + 1)(x + 1) = x^2 + 2x + 1 <----- perfect square trinomial

OpenStudy (anonymous):

ooh, okay, thanks

OpenStudy (cwrw238):

its called a trinomial because it has 3 terms

OpenStudy (cwrw238):

so can you answer the question?

OpenStudy (anonymous):

it's 1, right?

OpenStudy (cwrw238):

no , i'm talking about x^2 - x divide -1 by 2 gives -1/2 now square -1/2 (-1/2) * (-1/2) = ?